a balloon starts rising from the ground from rest at some constant acceleration. After some time, a stone is dropped from it. If the stone reaches the ground in the same time in which balloon reached the dropping point from ground , find the acceleration of the balloon.

Grade:11

1 Answers

Arun
3 years ago
1) Initially, 
Initial velocity, u = 0m/s 
acceleration, a =1.25 m/s^2 
Velocity of balloon after 8 s, v = u +at 
=> v = 0 + 1.25 * 8 = 10m/s 
2) Displacement of balloon in 8 s, 

3) After t = 8s, a stone is dropped. 
Initial velocity of stone at time of fall is same as that of balloon at that time :
That is velocity of stone,when it is dropped is 10m/s upward. 
Initial velocity, u = 10m/s
Displacement, s = -40m ( - means downward) 
Balloon is left. 
Acceleration, a = -10 m/s^2 

So, time taken by stone to reach the
t = 4 s, t = -2 (rejected) 
Therefore, Stone will take 4 seconds to reach the ground after being dropped from balloon at height h =40m from ground.